# Accelerated signal-to-noise ratio interpolation across the gravitational-wave signal manifold powered by artificial neural networks

### v1

@@ IV Conclusion and future work
In low-latency, the GstLAL search pipeline functions as a collection of microservices that loosely communicate with one another. Each microservice controls the filtering (among other tasks) for small groups of sub-banks. The microservices perform matched filtering in the reduced space described by the SVD bases, but combine these outputs to provide a physical SNR associated with the parameters of some template within the bank. This work shows that the intermediate SNRs collected by the basis vectors can be mapped to arbitrary points in the sub-bank in near-real-time. If a new microservice is created to ingest small snippets of intermediate SNRs from each filtering job, then the lightweight model described here can be used to rapidly estimate the SNRs across the entire search space. This effectively amortizes the costly likelihood calculation utilized by Bayesian inference based parameter estimation pipelines and provides the SNR time-series across the search space at arbitrary density.
This has the potential to both accelerate and improve the accuracy of near-real-time parameter estimation based tasks. Low-latency localization is performed using bayestar Singer and Price (2016), which fixes the masses and spins to the maximum search likelihood estimate and performs Bayesian inference over the extrinsic parameters. Though accurate, it is well known that full parameter estimation produces systematically more accurate skymaps, albeit at significantly greater computational cost. SNR interpolation would allow algorithms such as bayestar to marginalize over masses and spins to achieve a higher accuracy.
Source classifications using Bayesian inference are presently done with medium latency ($\mathcal{O}(\mathrm{minutes})$) algorithms that perform full parameter estimation Rose et al. (2022); Morisaki et al. (2023), but it is possible that these could be accelerated even more with our interpolant. Future work will discuss the connection of this technique to the RapidPE algorithm Rose et al. (2022); Pankow et al. (2015), which utilizes adaptive mesh refinement Berger and Colella (1989), in the pursuit of $\mathcal{O}(\mathrm{seconds})$ parameter estimation.

### v2

@@ IV Conclusion and future work
In low-latency, the GstLAL search pipeline functions as a collection of microservices that loosely communicate with one another. Each microservice controls the filtering (among other tasks) for small groups of sub-banks. The microservices perform matched filtering in the reduced space described by the SVD bases, but combine these outputs to provide a physical SNR associated with the parameters of some template within the bank. This work shows that the intermediate SNRs collected by the basis vectors can be mapped to arbitrary points in the sub-bank in near-real-time. If a new microservice is created to ingest small snippets of intermediate SNRs from each filtering job, then the lightweight model described here can be used to rapidly estimate the SNRs at arbitrary points in the search space. This effectively amortizes the costly likelihood calculation utilized by Bayesian inference based parameter estimation pipelines and provides the SNR time-series across the search space at arbitrary density.
This has the potential to both accelerate and improve the accuracy of near-real-time parameter estimation based tasks. Low-latency localization is performed using bayestar Singer and Price (2016), which fixes the masses and spins to the maximum search likelihood estimate and performs Bayesian inference over the extrinsic parameters. Though accurate, it is well known that full parameter estimation produces systematically more accurate skymaps, albeit at significantly greater computational cost. SNR interpolation would allow algorithms such as bayestar to marginalize over masses and spins to achieve a higher accuracy. Alternatively, it could be used to maximize the SNR via brute-force. After identifying candidates in low-latency, some detection pipelines employ an SNR optimization task to scan the local parameter space for slightly louder triggers. This interpolator could be used similarly, while probing a much denser space than current algorithms. While this will likely identify candidates with $\lesssim 1\%$ higher SNR, it will still lead to systematically tighter localizations.
Source classification using Bayesian inference is presently done with medium latency ($\mathcal{O}(\mathrm{minutes})$) algorithms that perform full parameter estimation Rose et al. (2022); Morisaki et al. (2023), but it is possible that these could be accelerated even more with our interpolant. These algorithms exploit the separability of aligned-spin gravitational wave signals to efficiently evaluate the GW likelihood for arbitrary extrinsic parameters. The method described here could similarly accelerate likelihood evaluations at arbitrary intrinsic parameters. Future work will discuss the connection of this technique to the RapidPE algorithm Rose et al. (2022); Pankow et al. (2015), which utilizes adaptive mesh refinement Berger and Colella (1989), in the pursuit of $\mathcal{O}(\mathrm{seconds})$ parameter estimation.